I need help with system of equations, I have a question on my homework it is, "A rectangle has a perimeter of 36 cm. Its width is 4 cm less than its length. Find the length and width of the rectangle."
I need to label the variables and create 2 equations, So far i have x=width y=length and 1 equation x=36-4.
I need help with the second equation. I hope this wasn't to long of a question. Thank you!

Answer by mananth(16946)

length = x
width = y
y=x-4..............................(1)
perimeter = 2(L+W)
2(x+y)= 36..........................(2)
substitute y in this equation
2(x+x-4)=36
2(2x-4)=36
/2
2x-4 = 18
2x=18+4
2x=22
/2
x=11 cm the length
width = 11-4 = 7cm